OpenStudy (anonymous):

please help me find h on this pythagoream problem: h^2=a^2+(a/2)^2 answer a.(a√5)/2 b.a√3 c.a√5 d.(a√3)/2

OpenStudy (mathmale):

\[h^2=a^2+(a/2)^2 = a ^{2}+\frac{ a ^{2} }{ 2^{2} }=a ^{2}+\frac{ a ^{2} }{ 4 }\]Simplify this expression.

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

is that like a sum of fraction right i just put 1 under a^2

OpenStudy (anonymous):

so the result is 5a^2/4

OpenStudy (mathmale):

Experiment first: Try adding 1/4 to 1 (or 1 to 1/4). Yes, your result is correct. Congrats!

OpenStudy (anonymous):

but i dont see that result on the options please help me/

OpenStudy (mathmale):

The quantity that we calculated was h^2. To find h, merely take the square root of our expression for h^2. Suggest choosing the positive square root.

OpenStudy (anonymous):

ok so h=√5a^2/√4

OpenStudy (anonymous):

or h=√5a^2/4